# Continuous-time persuasion by filtering

### v1

@@ 2.1 The informative value of smart meters
In what follows, we focus on the social economic value of smart meters in their role as information provision tools to consumers. We frame the problem in a persuasion framework in continuous-time. At initial time, a regulated electricity producer acting as a monopolist on the market considers investing in a device, the smart meter, designed to provide information to the consumers on their real-time consumption so that they can adjust their behaviour in response to fluctuating electricity prices. The device has several features that can be resumed by its ability to provide an accurate description of the continuous consumption of the consumer. Would the cost of such devices be negligible, the attention of the consumer infinite, and the electricity consistent with marginal costs of production, the solution is simply to install the most precise device and to provide full information to the consumer (see Aïd et al. (2023)). In a more realistic scenario, where the device is costly and consumers face frictions in changing their consumption habits, the producer has to make a trade-off choice between the rate at which information is provided to consumers and the induced investment cost.
@@
where $u_{0}>0$ and $c(v):=\frac{1}{2}v^{2}/\gamma$ is the cost of control and $\gamma>0$. The utility function of the consumer translates its preference for its average level of consumption $\ell$. Besides, the larger the parameter $u_{0}$, the larger the consumer gives importance to the induced comfort from this level of consumption. The real-time price of electricity is assumed to be an affine function of the consumption rate given by $p(x):=p_{0}+p_{1}x$, with $p_{0},p_{1}>0$. The consumer wishes to maximise the utility of electricity consumption minus friction costs and electricity bills. Namely, he solves the following stochastic control problem:
@@
where the admissible set of controls $\mathcal{A}^{\rm R}$ is specified in Section 3. At time $t\geq 0$, the consumer has access to a measure $M_{t}$ of his consumption $X_{t}$ thanks to the information device installed by the regulator, the smart meter. The measure of the consumption follows the dynamics
@@
where $b$ and $\sigma$ are non-negative parameters fixed by the regulator. The measure process $M$ is the information process generated by the smart meter. We also call $M$ the message process. The process $M$ captures two opposite situations: when $b\neq 0$ and $\sigma=0$, the consumer can estimate the true value of his consumption, which corresponds to the case of full information. On the other hand, when $b=0$ and $\sigma>0$, this corresponds to the absence of information provision. Between those two extreme situations lies a continuous range of communication device efficiency. The ratio ${b}/{\sigma}$ is called hereafter the precision of the device.
We denote by $h(b,\sigma)$ the investment cost for a device with characteristics $(b,\sigma)$. Such a function satisfies $h(0,\sigma)=0$, for all $\sigma>0$, and $h(b,0^{+}):=\lim_{\sigma\to 0^{+}}h(b,\sigma)=+\infty$. This means that it is costless to provide no information, and it is infinitely costly to provide the continuous-time information on the process $X$ to the consumer. Electricity is produced at a cost $g(x)$ of the form $g(x):=g_{0}x+\frac{1}{2}g_{1}x^{2}$, with $g_{0},g_{1}>0$. The problem of the producer that designs the device and sells electricity is given by
$\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}J^{\rm{S}}(b,\sigma;v):=\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}h(b,\sigma)+\limsup_{T\to\infty}\mathbb{E}\Bigg{[}\frac{1}{T}\int_{0}^{T}\Big{(}g(X_{t})-p(X_{t})X_{t}\Big{)}\mathrm{d}t\Bigg{]}.$ (2.5)
As in Aïd et al. (2023), this criterion of the producer could lead to an optimal solution that would consist in providing no information if the electricity price function $p$ is too large compared to the marginal cost of production $g^{\prime}$.
@@
$\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}\widetilde{J}^{\rm S}(b,\sigma):=$ $\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}h(b,\sigma)+\mathbb{E}\left[g(X^{*}_{\infty})-p(X^{*}_{\infty})X^{*}_{\infty}\right],$
@@
where
$\mathrm{d}X^{i}_{t}=-\kappa(X^{i}_{t}-\ell)\mathrm{d}t+v^{i}_{t}\mathrm{d}t+\mathrm{d}W^{i}_{t},\quad X_{0}^{i}=x_{0}^{i}>0,$ (2.7)
@@
$J^{\rm{S}}(b,\sigma):=h(b,\sigma)+\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}\Big{[}\int_{0}^{T}\Big{(}\frac{1}{N}\sum_{i=1}^{N}g(X^{i}_{t})-p(\bar{X}_{t})\bar{X}_{t}\Big{)}\mathrm{d}t\Big{]}.$ (2.9)
@@
where
$\mathrm{d}X_{t}=-\kappa(X_{t}-\ell)\mathrm{d}t+v_{t}\mathrm{d}t+\mathrm{d}W_{t},$ (2.11)
@@
$J^{\rm{S}}(b,\sigma):=h(b,\sigma)+\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}\Big{[}\int_{0}^{T}\Big{(}g(X_{t})-p(m)m\Big{)}\mathrm{d}t\Big{]}.$ (2.13)
@@ 2.2 Carbon footprint accounting rules
with $\lambda_{a},\lambda_{q}$ non-negative parameters and where $q\in\mathbb{R}$ is to be determined at equilibrium by the (mean-field) consistency condition that in this specific case reads as
@@
When $\epsilon=0$, the representative firm only wishes to be close to the average of the crowd of firm. On the other hand, when $\epsilon>0$, it means that the representative firm wishes to have a carbon footprint below the average, i.e., to be a best-in-class company in terms of carbon footprint reduction. The non-negative parameters $\lambda_{q}$ and $\lambda_{a}$ capture the firm’s prefernces between striving to be the best-in-class versus sticking to his own target carbon footprint, $a$, that can actually differ from the carbon footprint natural value of $\ell$. For a $\lambda_{q}$ close to $1$ and $\lambda_{a}$ near to zero, firms have a high tendency to lower their carbon footprint below the average level of their peers. This features captures preferences in the spirit of ranking games (see Bayraktar and Zhang (2016) and the references therein).
@@
$\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}h(b,\sigma)+\limsup_{T\to\infty}\frac{1}{T}\int_{0}^{T}\mathbb{E}\big{[}c(X_{t}^{*})^{2}\big{]}\mathrm{d}t,$ (2.18)
@@ 3.3 The Receivers’ ergodic mean-field game
Let us now consider a mean-field game (MFG) generalisation of the initial one-Receiver problem, as introduced at the beginning of Section 3, to address a case where we have a large population of identical Receivers getting information from the same Sender. MFGs have been pioneered by Lasry and Lions (2007) and Huang et al. (2006) as “natural” limit for large population stochastic differential games where the players are identical and they are linked by a mean-field interaction. In case of an ergodic criterion for all players, the latter means that each player interact with the long term empirical distribution of the population’s states and (possibly) actions. Therefore, when passing to the limit with respect to the number of players, it suffices to study the problem of a representative player interacting with a distribution over the state space, which at equilibrium is equal to the stationary distribution of the representative player’s state. When the objective functional is linear-quadratic as it is the case here, the distribution over the state space appears along the resolution process only through its stationary mean and (possibly) covariance. We refer to the two-volume book Carmona and Delarue (2019) for a full probabilistic treatment of MFGs and for a wide range of applications. Finally, among many papers on MFG with ergodic criterion, we refer to Bardi and Priuli (2013) and Bardi and Priuli (2014) for the analytic approach in the linear-quadratic case and to Cohen and Zell (2023), Cohen and Zell (2024) and Cao et al. (2023) for the probabilistic approach.
@@ 4.1 The informative value of smart meters
where we recall that the dynamics of the innovation process $dI_{t}$ is given by $dI_{t}=\mathrm{d}M_{t}-b\widehat{X}^{*}_{t}\mathrm{d}t=b(X^{*}_{t}-\widehat{X}_{t}^{*})\mathrm{d}t+\mathrm{d}B_{t}$, with $B$ and $W$ independent Brownian motions. Recall that, the volatility $\sigma$ of the information process $M$ is normalised to $1$.
@@
$\lim_{b\searrow 0}\sigma_{\infty}^{2}(b)=\sigma_{\infty}^{2}(0^{+})=\frac{1}{2\kappa}=:\bar{v},\quad\lim_{b\to+\infty}\sigma_{\infty}^{2}(b)=\sigma_{\infty}^{2}(+\infty)=\frac{1}{2\beta}=:\underline{v}.$
@@
$\inf_{b\geq 0}J^{\rm S}(b,\sigma):=$ $\inf_{b\geq 0}\frac{1}{2}\big{(}g_{1}-2p_{1}\big{)}\sigma_{\infty}^{2}(b)+h(b),$ (4.4)
where $h$ is the cost function of the device and where we restricted $b$ to be positive, as only its square intervenes in the optimisation problem of the Sender. Thus, as already exhibited in Aïd et al. (2023), it is necessary that the price of electricity is not greater than the marginal cost of production for the Sender to have an incentive to reduce the volatility. Note that, in the case where $g^{\prime}(x)=2g_{1}=p_{1}$, i.e., the electricity is priced at exactly the marginal cost of production, there is no benefit in developing an information device such as smart meters. In our model, only imperfections in the transmission to the consumers of certain production costs justify the development of an information solution. From now on, we assume that $0<g_{1}-2p_{1}$ $=$ $k_{\rm I}$, where $k_{\rm I}$ stands for the cost pricing imperfection.
@@
$\inf_{z\in(\underline{v},\bar{v}]}\frac{1}{2}k_{\rm I}z+H(z),$
where $z:=\sigma^{2}_{\infty}(b)$ and $H$ is such that $H(\bar{v})=0$ and $H(\underline{v})=+\infty$. This equivalent representation expresses the minimisation problem of the Sender directly using the cost of reducing the variance of the consumption. Thus, it is immediate that the optimal variance reduction should satisfy
@@
$H(z)=\frac{r\eta}{2}\Big{(}\frac{\Delta v}{z-\underline{v}}-1\Big{)},\quad z\in(\underline{v},\bar{v}],\quad\Delta v:=\bar{v}-\underline{v},$
@@
$z^{\ast}:=\begin{cases}\bar{v}&k_{\rm I}\Delta v\leq r\eta,\\ &\\ \underline{v}+\sqrt{\frac{\eta\Delta v}{k_{\rm I}}},&r\eta<k_{\rm I}\Delta v.\end{cases}$
Figure 1 (a) provides a representation of $\sigma^{2}_{\infty}(b)$ together with the optimal volatility value $z^{\ast}=\sigma_{\infty}^{2}(b^{\ast})$. The product $k_{\rm I}\Delta v$ captures two relevant terms in the decision process of the Sender: the higher it is, the higher the value of an information device. This quantity can be large either because of a significant imperfection of pricing cost $k_{\rm I}$ or because the potential variance reduction $\Delta v$ is large.
We turn now to the case of a mean-field of consumers in interaction though the electricity price. For a given value of $m$, the optimal control of the consumer is given by
@@
At equilibrium, the long-run consumption of a crowd of undistinguishable consumers is lower than the one of a representative agent. But, this result is an effect of the mean-field interaction not an effect of information provision. Furthermore, at equilibrium, the stationary distribution of the process $X^{*}$ is the Gaussian distribution $\mathcal{N}(m^{\ast}_{\infty},\sigma^{\ast}_{\infty}(b))$ with
@@
$\inf_{b\geq 0}J^{\rm S}(b,\sigma):=$ $\frac{1}{2}g_{1}\sigma_{\infty}^{2}(b)+h(b),$ (4.9)
because the income term $p(m^{\ast}_{\infty})m^{\ast}_{\infty}$ does not depend on $b$. Hence, the cost pricing imperfection increases from $k_{\rm I}=g_{1}-2p_{1}$ to a strictly greater value $k_{\rm I}^{\rm mf}:=g_{1}$. Appropriate volatility reduction follows with a higher threshold.
@@
Figure 1: (a) The variance $\sigma^{2}_{\infty}(b)$ of the stationary law of the consumption process $X$. Three tajectories of the consumption process $X$ and its posterior estimation $\widehat{X}$ for (b) $b=0$ (c) $b=5.5$ (d) $b=55$. Parameters: $\kappa=0.5$, $\ell=1$, $u_{0}=100$, $\gamma=1$, $p_{0}=50$, $p_{1}=100$; $g_{1}=400$, $h=0.5$.

### v2

@@ 2.1 The informative value of smart meters
In what follows, we focus on the social economic value of smart meters in their role as information provision tools to consumers. We frame the problem in a persuasion framework in continuous-time. At initial time, a regulated electricity producer acting as a monopolist on the market considers investing in a device, the smart meter, designed to provide information to the consumers on their real-time consumption so that they can adjust their behaviour in response to fluctuating electricity prices. The device has several features that can be resumed by its ability to provide an accurate description of the continuous consumption of the consumer. This accuracy can be thought as the frequency at which the device provides information on the consumption of the consumers. Indeed, prior to the development of smart meters, the electricity consumption of households were measured once a year, and bills that were sent very month or every two months, were based on past estimation of the consumption. Now, with smart meters, it is possible to measure the electricity consumption at a daily basis, or an hourly basis and even at a 10-minutes basis. The consumers could reveice a bill every hour. Would the cost of such devices be negligible, the attention of the consumer infinite, and the electricity consistent with marginal costs of production, the solution is simply to install the most precise device and to provide full information to the consumer (see Aïd et al. (2023)). However, it is quite clear that measuring and delivering to the consumers its consumption at a one-minute basis should be (much) more costly than a simple annual measure. Thus, when the device is costly and consumers face frictions in changing their consumption habits, the producer has to make a trade-off choice between the rate at which information is provided to consumers and the induced investment cost.
@@
where $u_{0}>0$ and $c(v):=\frac{1}{2}v^{2}/\gamma$ is the cost of control and $\gamma>0$. The utility function of the consumer translates its preference for its average level of consumption $\ell$. Besides, the larger the parameter $u_{0}$, the larger the consumer gives importance to the induced comfort from this level of consumption. The real-time price of electricity is assumed to be an affine function of the consumption rate given by $p(x):=p_{0}+p_{1}x$, with $p_{0},p_{1}>0$. This price can be seen as the spot price of electricity. The consumer wishes to maximise the utility of electricity consumption minus friction costs and electricity bills. Namely, he solves the following stochastic control problem:
@@
where the admissible set of controls $\mathcal{A}^{\rm R}$ is specified in Section 3. Note that the model assumes that the instantaneous price $p(X_{t})$ to be non-observable, i.e. the consumer does not observe the spot price of electricity. This translates the idea described above that the consumption and its cost are known to the consumer at a frequency which depends on the metering device. Indeed, at time $t\geq 0$, the consumer has access to a measure $M_{t}$ of his consumption $X_{t}$ thanks to the information device installed by the producer. The measure of the consumption follows the dynamics
@@
where $b$ and $\sigma$ are non-negative parameters characterising the smart meter. The measure process $M$ is the information process generated by the smart meter. When $b=0$, the consumers is left with the only information of its average consumption. This situation can be interpreted as the pre-smart metering area when consumption was measured once a year. As $b$ increases and $\sigma$ remains constant, the consumer can infer more an more rapidely its true consumption compared to what he thinks it is. The ratio ${b}/{\sigma}$ is called hereafter the precision of the device (signal over noise ratio). Moreover, because only the ratio $b/\sigma$ matters in the design of the smart meter, we will normalise it to $1$, and reduce our smart meter design to $b$. We denote by $h(b)$ the investment cost for a device with characteristics $b$. Such a function satisfies $h(0)=0$, and $\lim_{b\to\infty}h(b)=+\infty$, and $h$ strictly convex. This means that it is costless to provide no information, and it is infinitely costly to provide the continuous-time information on the process $X$ to the consumer.
On the production side, electricity is produced at a cost $g(x)$ of the form $g(x):=g_{0}x+g_{1}x^{2}$, with $g_{0},g_{1}>0$. The problem of the producer that designs the device and sells electricity is given by
$\inf_{b\in\mathcal{A}^{\rm S}}J^{\rm{S}}(b;v):=\inf_{b\in\mathcal{A}^{\rm S}}h(b)+\limsup_{T\to\infty}\mathbb{E}\Bigg{[}\frac{1}{T}\int_{0}^{T}\Big{(}g(X_{t})-p(X_{t})X_{t}\Big{)}\mathrm{d}t\Bigg{]}.$ (2.5)
@@
$\inf_{b\in\mathcal{A}^{\rm S}}\widetilde{J}^{\rm S}(b):=$ $\inf_{b\in\mathcal{A}^{\rm S}}h(b)+\mathbb{E}\left[g(X^{*}_{\infty})-p(X^{*}_{\infty})X^{*}_{\infty}\right],$
@@
where $\mathrm{d}X^{i}_{t}=-\kappa(X^{i}_{t}-\ell)\mathrm{d}t+v^{i}_{t}\mathrm{d}t+\mathrm{d}W^{i}_{t},\quad X_{0}^{i}=x_{0}^{i}>0,$ (2.7)
@@
$J^{\rm{S}}(b):=h(b)+\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}\Big{[}\int_{0}^{T}\Big{(}\frac{1}{N}\sum_{i=1}^{N}g(X^{i}_{t})-p(\bar{X}_{t})\bar{X}_{t}\Big{)}\mathrm{d}t\Big{]}.$ (2.9)
@@
where $\mathrm{d}X_{t}=-\kappa(X_{t}-\ell)\mathrm{d}t+v_{t}\mathrm{d}t+\mathrm{d}W_{t},$ (2.11)
@@
$J^{\rm{S}}(b):=h(b)+\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}\Big{[}\int_{0}^{T}\Big{(}g(X_{t})-p(m)m\Big{)}\mathrm{d}t\Big{]}.$ (2.13)
@@ 2.2 Carbon footprint accounting rules
with $\gamma>0$, $c(v):=\frac{1}{2}\frac{v^{2}}{\gamma}$ the cost of control and $\lambda_{a},\lambda_{q}$ non-negative parameters and where $q\in\mathbb{R}$ is to be determined at equilibrium by the (mean-field) consistency condition:
@@
When $\epsilon=0$, the representative firm only wishes to be close to the average of the crowd of firm. On the other hand, when $\epsilon>0$, it means that the representative firm wishes to have a carbon footprint below the average, i.e., to be a best-in-class company in terms of carbon footprint reduction. The non-negative parameters $\lambda_{q}$ and $\lambda_{a}$ capture the firm’s preferences between striving to be the best-in-class versus sticking to his own target carbon footprint, $a$, that can actually differ from the carbon footprint natural value of $\ell$. For a $\lambda_{q}$ close to $1$ and $\lambda_{a}$ near to zero, firms have a high tendency to lower their carbon footprint below the average level of their peers. This features captures preferences in the spirit of ranking games (see Bayraktar and Zhang (2016) and the references therein).
@@
$\inf_{(b,\sigma)\in\mathcal{A}^{\rm S}}h(b,\sigma)+\limsup_{T\to\infty}\frac{1}{T}\int_{0}^{T}\frac{1}{2}\mathbb{E}\big{[}c(X_{t}^{*})^{2}\big{]}\mathrm{d}t,$ (2.18)
@@ 3.3 The Receivers’ ergodic mean-field game
Let us now consider a mean-field game (MFG) generalisation of the initial one-Receiver problem, as introduced at the beginning of Section 3, to address a case where we have a large population of identical Receivers getting information from the same Sender. MFGs have been pioneered by Lasry and Lions (2007) and Huang et al. (2006) as “natural” limit for large population stochastic differential games where the players are identical and they are linked by a mean-field interaction. In case of an ergodic criterion for all players, the latter means that each player interacts with the long term empirical distribution of the population’s states and (possibly) actions. Therefore, when passing to the limit with respect to the number of players, it suffices to study the problem of a representative player interacting with a distribution over the state space, which at equilibrium is equal to the stationary distribution of the representative player’s state. When the objective functional is linear-quadratic as it is the case here, the distribution over the state space appears along the resolution process only through its stationary mean and (possibly) covariance. We refer to the two-volume book Carmona and Delarue (2019) for a full probabilistic treatment of MFGs and for a wide range of applications. Finally, among many papers on MFG with ergodic criterion, we refer to Bardi and Priuli (2013) and Bardi and Priuli (2014) for the analytic approach in the linear-quadratic case and to Cohen and Zell (2023), Cohen and Zell (2024) and Cao et al. (2023) for the probabilistic approach.
@@ 4.1 The informative value of smart meters
where we recall that the dynamics of the innovation process $dI_{t}$ is given by $dI_{t}=\mathrm{d}M_{t}-b\widehat{X}^{*}_{t}\mathrm{d}t=b(X^{*}_{t}-\widehat{X}_{t}^{*})\mathrm{d}t+\mathrm{d}B_{t}$, with $B$ and $W$ independent Brownian motions. Recall that, the volatility $\sigma$ of the information process $M$ has been normalised to $1$.
@@
$\lim_{b\searrow 0}\sigma_{\infty}^{2}(b)=\sigma_{\infty}^{2}(0^{+})=\frac{1}{2\kappa}=:\bar{\sigma}^{2},\quad\lim_{b\to+\infty}\sigma_{\infty}^{2}(b)=\sigma_{\infty}^{2}(+\infty)=\frac{1}{2\beta}=:\underline{\sigma}^{2}.$ (4.4)
@@
$\inf_{b\geq 0}J^{\rm S}(b):=$ $\inf_{b\geq 0}\big{(}g_{1}-p_{1}\big{)}\sigma_{\infty}^{2}(b)+h(b),$ (4.5)
where $h$ is the cost function of the device and where we restricted $b$ to be positive, as only its square intervenes in the optimisation problem of the Sender. Thus, as already exhibited in Aïd et al. (2023), it is necessary that the price of electricity is not greater than the marginal cost of production for the Sender to have an incentive to reduce the volatility. Note that, in the case where $g^{\prime}(x)=2g_{1}x=2p_{1}x$, i.e., the electricity is priced at exactly the marginal cost of production, there is no benefit in developing an information device such as smart meters. In our model, only imperfections in the transmission to the consumers of certain production costs justify the development of an information solution. From now on, we assume that $0<g_{1}-p_{1}$ $=$ $k_{\rm I}$, where $k_{\rm I}$ stands for the cost pricing imperfection.
@@
$\inf_{z\in(\underline{\sigma}^{2},\bar{\sigma}^{2}]}\frac{1}{2}k_{\rm I}z+H(z),$
where $z:=\sigma^{2}_{\infty}(b)$ and $H$ is such that $H(\bar{\sigma}^{2})=0$ and $H(\underline{\sigma}^{2})=+\infty$. This equivalent representation expresses the minimisation problem of the Sender directly using the cost of reducing the variance of the consumption. Thus, it is immediate that the optimal variance reduction should satisfy
@@
$H(z)=\frac{r\eta}{2}\Big{(}\frac{\Delta\sigma^{2}}{z-\underline{\sigma}^{2}}-1\Big{)},\quad z\in(\underline{\sigma}^{2},\bar{\sigma}^{2}],\quad\Delta\sigma^{2}:=\bar{\sigma}^{2}-\underline{\sigma}^{2},$
@@
$(\sigma^{\ast})^{2}:=\begin{cases}\bar{\sigma}^{2}&k_{\rm I}\Delta\sigma^{2}\leq r\eta,\\ &\\ \underline{\sigma}^{2}+\sqrt{\frac{r\eta\Delta\sigma^{2}}{k_{\rm I}}},&r\eta<k_{\rm I}\Delta\sigma^{2}.\end{cases}$
Figure 1 (a) provides a representation of $\sigma^{2}_{\infty}(b)$ together with the optimal volatility value $\sigma^{\ast}=\sigma_{\infty}(b^{\ast})$. The product $k_{\rm I}\Delta\sigma^{2}$ captures two relevant terms in the decision process of the Sender: the higher it is, the higher the value of an information device. This quantity can be large either because of a significant imperfection of pricing cost $k_{\rm I}$ or because the potential variance reduction $\Delta\sigma^{2}$ is large.
We turn now to the case of a mean-field of consumers in interaction through the electricity price. For a given value of $m$, the optimal control of the consumer is given by
@@
At equilibrium, the long-run consumption of a crowd of undistinguishable consumers is lower than the one of a representative agent. But, this result is an effect of the mean-field interaction not an effect of information provision. Furthermore, at equilibrium, the stationary distribution of the process $X^{*}$ is the Gaussian distribution $\mathcal{N}(m^{\ast}_{\infty},(\sigma^{\ast}_{\infty})^{2}(b))$ with
@@
$\inf_{b\geq 0}J^{\rm S}(b,\sigma):=$ $g_{1}\sigma_{\infty}^{2}(b)+h(b),$ (4.10)
because the income term $p(m^{\ast}_{\infty})m^{\ast}_{\infty}$ does not depend on $b$. Hence, the cost pricing imperfection increases from $k_{\rm I}=g_{1}-p_{1}$ to a strictly greater value $k_{\rm I}^{\rm mf}:=g_{1}$. Appropriate volatility reduction follows with a higher threshold.
@@
Figure 1: (a) The variance $\sigma^{2}_{\infty}(b)$ of the stationary law of the consumption process $X$. Three trajectories of the consumption process $X$ and its posterior estimation $\widehat{X}$ for (b) $b=0$ (c) $b=5.5$ (d) $b=55$. Parameters: $\kappa=0.5$, $\ell=1$, $u_{0}=100$, $\gamma=1$, $p_{0}=50$, $p_{1}=100$; $g_{1}=400$, $h=0.5$.